OpenStudy (anonymous):

The diagonals of a rhombus are 8 and 10. The area of the rhombus is

OpenStudy (anonymous):

a-20 b-40 c-80

OpenStudy (goldphenoix):

The formula to finding the area of the rhombus is: \[\large \large \frac{ D1*D2 }{ 2 } = Area\] D1 = Diagonal #1 D2 = Diagonal #2 (The other diagonal) So by plugging in what we know, our equation should look like: \[\large \large \frac{ 8*10 }{ 2 }=Area\] ^What would the area be?

OpenStudy (anonymous):

40?

OpenStudy (goldphenoix):

Correct!